OpenStudy (anonymous):

What is the common difference in the sequence 8, 16, 24, 32, . . .? A. 24 B. 16 C. –8 D. 8

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Thanks! (:

OpenStudy (anonymous):

16 - 8 = 24 - 16 = 32 - 24 = 8 (D) @Alicia88 Solutions >> Answers

OpenStudy (anonymous):

@jenniferlynn92 What does "difference" mean? What operation is it?